Reports 1-1 of 1 Clear search Modify search
AOS (Baffles & Beam dumps)
tomotada.akutsu - 23:40 Thursday 20 December 2018 (7469) Print this report
Evacuating the EXT chamber

K. Tanaka, Akutsu

Started the evacuation of the EXT chamber this afternoon around 15:15 (Fig 1). My concern was a hard storm in the chamber during the pumping; once that would happen, that would make damages to fine wires for coils in the chamber. So, I tried to tune the conductance of the evacuated air flow; firstly opened the angle valve very slightly and then turned on the roots pump along with monitoring all the six LVDT outputs and the CCD images for IR and green; note that they were not necessarily resonated. Actually, to share the limited time properly the other commissioning works were ongoing, but their randomly-flashing images were sufficiently useful to know how they were drifting during the evacuation. Then I increased the opening aperture of the angle valve by judging from the several sensors like the LVDTs and the CCDs; I'm sorry but it is hard to explain...shoujin aru nomi! or don't think; feel.

The LVDTs showed the TMS-VIS's suspended table was drifting especially in the vertical direction (Fig 2). This could correspond to losing the buoyance force by air, but need to be confirmed quantitatively later; in fact, the suspended table was sinking along with loweinrg the pressure, and that is consistent with the buoyance assumption qualitatively. My concern was that the IR or green spots would be drifting out of the CCD views, but the drifts were not such largelaugh!  and the spots were still within the CCDs' views. Probably, later, we need to tweak the steering mirrors in front of the CCDs so that they should come to their centers.

During the pumping down, there was an earthquake at around 16:00; actually I was surprised that I was able to "feel" the earthquake in the tunnel! This was the first time for me to feel an earthquake in the mine... The shock was like an impulse, made me move in the perpendicular direction with respect to the X arm (I was sitting on a chair, and felt the chair moved suddenly!). I fisrtly felt that someone kicked my ass... you ha shook! Probably my sensitivity was high as I was sitting on a chair. At the same time I think I heard a explosive low-frequency sound... so firstly I thought anywhere the vacuum chamber got implosion or explosion such as happened in the iKAGRA era!! Anyway, every suspension in KAGRA would get shook and the result was that I lost any light images on the CCDs, but we continued the evacuation of the EXT.

Of course, the TMS-VIS was also shook due to the earthquake (Fig 3). After damped (Fig 4), the LVDTs came back to the line of the original trend; so nice...

Then at 16:30, the pressure reached 15 Pa (Fig 5), so I turned on the turbo molecular pump (TMP) and then the pressure was immediately getting down to the order of 0.1Pa, so we left them and went back to Mozumi. Let's see tomorrow.

Note: how to read the indication of Fig 1 or 5; the barometer is a CC-10, a combination gauge to monitor from 1e-7 to 1e5 Pa. For example, in Fig 1, it says 9.9e4 Pa.

Images attached to this report
Comments to this report:
yutaro.enomoto - 18:43 Friday 21 December 2018 (7476) Print this report

= Minor adjustment of the alignment for green and Opening test of the GV between EXC and EXT =

One day after starting the evacuation, the GV between EXC and EXT was opened for a while.
No significant alignment change was observed, but the main laser power measured by the PD increased by ~5 %.

- Alignment adjust for GigE camera and DC PD for green.
Since the green beam was off from the center of the camera or the PD by 1-2 mm, I adjusted the alignment for them.
I used the steering mirror (or BS) just before each detector.
The first and second picture show the camera image before and after the adjust. 

- Open (and close) of the GV
16:50 JST, I opened the GV between EXC and EXT, after talking with Saitoh-san.
The pressure at EXT was 4.6e-3 Pa before the open and 5.4e-4 Pa after the open.
Before and after the open, there were no significant changes in beam alignment of IR or green;
I put alignment references as third and fourth picture show, and the beam spot changes at the references were less than 1 mm. In fact I could not see any changes.
There were no changes found in the GigE images as well.
As for the transmission power, there was a change. See the fifth attachment.
It seems that the transmission power of IR increased by ~5%. This was a significant change.
(This was not due to alignment fluctuation, or increase of the input power.)

17:09 JST, I closed the GV.

Images attached to this comment
Search Help